# Complex-Valued Holographic Radiance Fields

### v1

@@ Abstract.
Modeling the full properties of light, including both amplitude and phase, in 3D representations is crucial for advancing physically plausible rendering, particularly in holographic displays. To support these features, we propose a novel representation that optimizes 3D scenes without relying on intensity-based intermediaries. We reformulate 3D Gaussian splatting with complex-valued Gaussian primitives, expanding support for rendering with light waves. By leveraging RGBD multi-view images, our method directly optimizes complex-valued Gaussians as a 3D holographic scene representation. This eliminates the need for computationally expensive hologram re-optimization. Compared with state-of-the-art methods, our method achieves 30x-10,000x speed improvements while maintaining on-par image quality, representing a first step towards geometrically aligned, physically plausible holographic scene representations.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXsubmissionid: Reviewccs: Computing methodologies Mixed / augmented realityccs: Computing methodologies 3D imagingccs: Computing methodologies Computer graphicsccs: Computing methodologies Rendering
Figure 1. Complex-Valued Holographic Radiance Fields. Instead of considering interference and diffraction as a secondary derivation from intensity data, our method explicitly treats the complex-valued holographic radiance fields as the primary optimization variable at the beginning of the scene modeling, jointly optimizing the 3D scene across multiple viewpoints to model intensity, interference, and diffraction of light in scene representation. The complex-valued 3D Gaussians will be projected to multiple depth planes, resulting in a complex 3D hologram via holographic rendering. This enables holographic reconstruction with depth-of-field effects: green boxes highlight the back focus region, and red boxes show the front focus region.
@@ 1. Introduction
Light exhibits numerous physical phenomena beyond intensity, such as spectral variation, polarization, diffraction, and interference (Goodman, 2005). Recent advancements in  (Kerbl et al., 2023) have enabled efficient novel-view synthesis for geometry-aware 3D scene representation. However, these approaches focus solely on intensity-based modeling and thus neglect the wave-optics properties introduced above. Modeling these wave-optics properties accurately is essential for physically plausible rendering of visual scenes. Among these properties, diffraction and interference are fundamental to holographic displays. Unlike conventional displays, holographic displays are an emerging technology that creates immersive 3D experiences. By reconstructing light diffraction—in other words, the phase from a hologram—holographic displays can reproduce the depth cues that our visual system perceives in natural environments (Kim et al., 2024).
is the computational backbone of holographic displays, enabling the synthesis of holograms that reconstruct 3D wavefronts. Existing methods decouple phase retrieval from scene representation, treating hologram synthesis as a post-processing step applied after the intensity distribution is determined. These methods typically estimate phase patterns that reproduce desired reconstructed intensity image for a fixed perspective (Shi et al., 2021, 2022; Choi et al., 2025; Kavaklı et al., 2023b; Chakravarthula et al., 2020; Choi et al., 2021a; Kavaklı et al., 2023a; Peng et al., 2020; Liu et al., 2023; Schiffers et al., 2023; Hossein Eybposh et al., 2020) and a narrow viewing angles (Chao et al., 2024; Wang et al., 2024; Chakravarthula et al., 2022; Kim et al., 2024; Chen et al., 2025). This viewpoint-specific wave propagation requires recalculating diffraction patterns when the geometric relationship between the object and hologram plane changes. In this context, phase values function as computational variables optimized for specific views, not as intrinsic properties of the 3D scene. As shown in Fig. 2, the estimated phase fails to align with the scene geometry across novel views, leading to visual artifacts during geometric transitions and physically inconsistent wave optics effects. Existing methods (Shi et al., 2022; Kavaklı et al., 2023b; Choi et al., 2025) typically require view-dependent transformations or complete recalculation of holograms for each viewpoint to maintain consistency.
Figure 2. Left: Traditional methods calculate hologram for a single view, but the wavefront collapses when the geometric configuration changes, requiring recalculation for each viewpoint. Right: A complex-valued holographic radiance field stores a scene-wide, geometrically aligned phase; the rainbow bands depict geometry-aligned phase.
@@
Our work aims to coherently merge traditional intensity-based radiance fields and holography. We present a novel framework that extends by using complex-valued Gaussian primitives to model both amplitude and phase information of the scene. Our method explicitly treats the complex-valued holographic radiance field as the primary optimization target at the beginning of scene modeling, without the need for per-view transformation. We jointly optimize the scene across multiple viewpoints to model interference and diffraction phenomena directly. We propose a differentiable rasterizer that effectively renders complex-valued Gaussian primitives using learnable probabilistic assignments. Unlike existing Eulerian-based methods that require per-primitive calculations scaling quadratically with the number of primitives, our Lagrangian-based approach is more effective due to the linear relationship between the number of primitives and complexity, achieving 30x-10,000x speed improvements over state-of-the-art methods (Choi et al., 2025). By optimizing directly in the complex domain across multiple views, our method achieves geometrically aligned, view-dependent diffraction and interference effects without relying on intensity-based intermediaries, representing a first step towards physically plausible holographic scene representations. Specifically, our contributions include:
  • We present complex-valued holographic radiance fields, a Lagrangian-based complex-valued Gaussian representation for transformation-free, geometrically aligned holographic rendering.
@@
  • We propose a novel differentiable rasterizer, enabling efficient rendering and optimization of complex-valued holographic radiance fields.
@@ 2. Related Work
Our approach bridges recent advances in neural rendering with holography, enabling the efficient generation of holograms from novel views. We will first review novel-view synthesis and then examine existing wave optics rendering techniques that also attempt to simulate light phenomena beyond intensity. Finally, we discuss the traditional methods in general, including point-based, polygon-based, and layer-based methods.
@@ 2.1. Novel-View Synthesis
Traditional novel-view synthesis methods focus exclusively on modeling intensity and appearance, which limits their applicability to emerging display technologies that require holographic properties. Early image-based rendering approaches (Chaurasia et al., 2013) and volumetric methods (Hedman et al., 2017; Riegler and Koltun, 2020) established the foundation for view-dependent appearance modeling. The introduction of  (Mildenhall et al., 2021) changed the field by learning continuous radiance functions that encode view-dependent color and opacity. Subsequent extensions improved quality and efficiency (Barron et al., 2021, 2022, 2023; Chen et al., 2022; Rebain et al., 2021; Niemeyer et al., 2024; Duckworth et al., 2024).
Point-based representations have recently advanced with  (Kerbl et al., 2023), emerging as a promising alternative to . offers a promising foundation for our work, as it encodes both geometry and appearance using oriented Gaussian primitives. Unlike implicit representations, provides explicit geometric primitives that naturally align with holographic modeling principles, but it is fundamentally limited to intensity-based representations. Our work extends to include intrinsic amplitude and phase properties, enabling the modeling of complex-valued holographic radiance fields that capture not just intensity, but also the interference and diffraction phenomena.
@@
The work by Kim  et al. (2023) extends beyond intensity-based representations to capture spectral and polarimetric properties of light, revealing richer material and structural characteristics of 3D scenes. Notably, spectral and polarization data can be supervised using ground-truth measurements obtained from specialized imaging devices. On the other hand, interference and diffraction phenomena arise from the accumulation of phase and amplitude values. Interference and diffraction effects lack direct measurement data and must be inferred through computational means, as holographic cameras capable of capturing both amplitude and phase information are not yet commodity equipment (An et al., 2020). Moreover, supporting simulations of holography in rendering pipelines is a common theme in research. For instance, Steinberg  et al. (2022; 2024) introduced wave optics formulations for path tracing, enabling partial coherence and diffraction effects. Choi  et al. (2025) proposed that transforms 2D Gaussian scenes into holograms by deriving closed-form solutions for angular spectrum. Chen  et al. (2025) proposed a two-stage method that combines pretrained intensity-based with U-Net (Ronneberger et al., 2015), enabling efficient hologram synthesis for novel views.
@@ 2.2. Eulerian-based Computer-Generated Holography
Holographic displays rely on , which algorithmically synthesizes diffraction and interference patterns to reconstruct 3D scenes with near-accurate depth cues. Conceptually, traditional methods follow a per-view Eulerian approach to wave propagation, where the coordinates on the hologram plane remain fixed, while the complex field varies when the relationship between the scene geometry and the viewpoint changes, requiring recalculation. The computed holographic patterns are implemented by , which modulates the phase and optionally amplitude of incident coherent beams to generate designed interference fringes. Through numerical wave propagation simulations (Kavaklı et al., 2022; Peng et al., 2020), generates holograms leading to high spatio-angular resolution, as an emerging screen technology. The accuracy and computational complexity of strongly depend on primitive representations, generally categorized into point-based, polygon-based, and layer-based methods.
Point-based represents scenes using discrete points emitting spherical waves. The wave field at the hologram, $P$, is calculated by summing waves from N points:
@@
where $A_{n}$ denotes amplitude, $k$ is the wave number, and $r_{n}$ is the propagation distance from point $n$ to $P(x,y)$. Reflecting the Huygens-Fresnel principle, Eq. (1) is highly flexible and physically accurate. Maimone  et al. (2017) presented point-source based computation for holographic near-eye displays for virtual and augmented reality. Shi  et al. (2021) leveraged point clouds and triangle-ray intersection for occlusion handling in large-scale hologram datasets. However, computational cost scales drastically with point density, and occlusion handling requires additional geometric computations.
@@
Polygon-based balances efficiency and accuracy by representing surfaces using triangular or polygonal facets emitting coherent wave fields. This method naturally integrates shading models and intrinsic occlusion handling through visibility algorithms (e.g. back-face culling and affine transformations) (Zhang et al., 2018b; Matsushima and Nakahara, 2009). Mesh-based includes topology information for precise occlusion handling (Ahrenberg et al., 2008; Yeom et al., 2022). While achieving realistic shading by combining diffraction with rendering pipelines, these methods suffer from increased algorithmic complexity and struggle with finely detailed or non-polygonal structures.
@@
Layer-based subdivides scenes into discrete depth planes, propagating complex fields from $P$ independently to each plane using operators like  (Matsushima and Shimobaba, 2009; Zhang et al., 2020; Zheng et al., 2024). The $P(x,y)$ is optimized so propagated results across all planes approximate the desired 3D scene. This method greatly accelerates computations through -based propagation, making it suitable for faster rendering and widely adopted in learned methods (Shi et al., 2022; Choi et al., 2021b; Peng et al., 2020; Zhan et al., 2024; Liu et al., 2023).
@@
While recent advancements enable realistic defocus effects and continuous pupil accommodation (Kim et al., 2024; Kavaklı et al., 2023a), existing light field–based 3D and 4D methods are typically optimized for narrow viewing directions, with computed phase patterns losing validity under significant geometric changes. We propose integrating strengths of both point-based and layer-based methods, utilizing 3D Gaussians as point-based primitives to encode scene geometry and radiance. Unlike existing Eulerian approaches, our method follows a Lagrangian-based formulation, treating amplitude and phase of light as intrinsic variables to Gaussian primitives. Compared with point clouds and polygon-meshes, Gaussians have learnable scale and rotation parameters, allowing them to capture anisotropic structures and fine-grained spatial variations effectively. This expressiveness supports both appearance modeling and geometrically faithful wave emitter distribution. Table 1 demonstrates our proposed method, enabling fast inference without per-view recalculation, natural defocus blur, on-par image quality, and geometrically aligned phase across novel views.
@@
GWS (2025) Moderate Yes Complex Good No No
@@
Table 1. Comparison of Gaussian primitives based hologram synthesis method. Inference refers to inference speed. Re refers to whether the method requires recalculation when geometric relationship between scene and observer changes. Type refers to Gaussian representation type. Quality refers to the image quality of reconstructed images from the holograms. Align refers to geometrically aligned phase across novel views. NDB refers to natural defocus blur.
Figure 3. Without the need for querying from an intensity-based radiance field, our approach models a complex-valued holographic radiance field using Gaussian primitives with intrinsic phase properties, enabling geometrically aligned phase relationships across viewpoints and efficient rendering through a differentiable multi-layer propagation pipeline.

### v2

@@ Abstract.
Modeling wave properties of light is an important milestone for advancing physically-based rendering. In this paper, we propose complex-valued holographic radiance fields, a method that optimizes scenes without relying on intensity-based intermediaries. By leveraging multi-view images, our method directly optimizes a scene representation using complex-valued Gaussian primitives representing amplitude and phase values aligned with the scene geometry. Our approach eliminates the need for computationally expensive holographic rendering that typically utilizes a single view of a given scene. This accelerates holographic rendering speed by 30x-10,000x while achieving on-par image quality with state-of-the-art holography methods, representing a promising step towards bridging the representation gap between modeling wave properties of light and 3D geometry of scenes.
@@
journalyear: 2024doi: XXXXXXX.XXXXXXXsubmissionid: Reviewccs: Computing methodologies Mixed / augmented realityccs: Computing methodologies 3D imagingccs: Computing methodologies Computer graphicsccs: Computing methodologies Rendering
Figure 1. Complex-Valued Holographic Radiance Fields. Our method explicitly treats the complex-valued holographic radiance fields as the primary optimization target, jointly optimizing the 3D scene across multiple viewpoints to model intensity, interference, and diffraction of light in scene representations. The complex-valued 3D Gaussians will be projected to multiple depth planes, resulting in a complex 3D hologram via holographic rendering. This enables holographic reconstruction with depth-of-field effects: blue boxes highlight the back focus region, and red boxes show the front focus region.
@@ 1. Introduction
Light exhibits numerous physical phenomena beyond mere intensity, including wave-optics properties such as spectral variations (Kim et al., 2023), polarization (Zhao et al., 2022), diffraction, and interference (Goodman, 2005). Accounting for wave-optics properties of light more accurately could help increase the visual fidelity in 3D scene representations (Wen et al., 2025). Most recently, 3D Gaussian splatting (3DGS) (Kerbl et al., 2023) has emerged as an effective 3D scene representation to synthesize novel-view intensity images. However, 3DGS neglects wave-optics properties of light. In addition, the applications of physically-based rendering and emerging holographic 3D displays (Kim et al., 2024) in computer graphics demand modeling these wave-optics properties to be accounted for.
Computer-Generated Holography (CGH) synthesizes holograms that faithfully reconstruct wavefronts emitted from 3D scenes. While CGH has been predominantly utilized for holographic display applications, it has recently garnered increasing interest from the rendering community for generating physically-plausible scenes in computer graphics (Steinberg et al., 2022, 2024). Existing CGH methods treat hologram synthesis as a post-processing task applied after the conventional rendering pipeline. These methods typically estimate holograms that reproduce the desired reconstructed intensity image for a fixed viewpoint (Shi et al., 2021) or a limited viewing angle (Chakravarthula et al., 2022; Chen et al., 2025). When such methods are queried with a new viewpoint, they have to be recalculated for a new hologram using only the provided viewpoint information. Most recently, the work by Choi  et al. (2025) generates holograms of 3D scenes represented with real-valued, intensity-based 2D Gaussian primitives. Their method introduces multi-view based approaches tailored for CGH for the first time, requiring an additional step of transformation for each queried viewpoint, greatly increasing computational complexity. These methods adopt an Eulerian paradigm, in which the hologram plane remains fixed while the complex field is recomputed at each spatial coordinate. In this context, previous CGH methods lack persistent geometric representations, computing holograms at the camera plane rather than coupling them to 3D scene geometry. Thus, such holographic scene representations without any transformation or recalculation would lead to noise-like visuals rather than perspective images when the viewpoint changes, as demonstrated in Fig. 2.
Figure 2. Left: Existing CGH methods do not preserve geometry of scenes under viewpoint changes, necessitating transformation or recalculation per-view. Right: Our complex-valued holographic radiance field offers a 3D consistent representation in contrast.
@@
Our work aims to merge 3DGS radiance fields and holographic rendering under one umbrella using complex numbers, which are widely adopted in holography for representing both amplitude and phase information from 3D scenes (Goodman, 2005). In this paper, we firstly introduce complex-valued Gaussian primitives. Our scene representation framework utilizes these complex-valued Gaussian primitives to model both amplitude and phase emitted from 3D scenes. Unlike previous intensity-based modeling methods (Choi et al., 2025), our method explicitly treats complex-valued holographic radiance fields as the primary optimization target, without the need for per-view transformation (Choi et al., 2025) or fixed viewpoint hologram recalculation (Chen et al., 2025). Secondly, in our framework the simulation of diffraction and interference phenomena is formulated as a differentiable rasterizer, rendering complex-valued 3D Gaussian primitives from novel views. In contrast to the existing CGH methods (Choi et al., 2025) that scale quadratically with the number of primitives, our approach achieves a linear relationship between the primitives and complexity, achieving 30x-10,000x speed improvements. By optimizing directly in the complex-valued domain across multiple viewpoints, our method achieves view-dependent diffraction and interference effects without relying on intensity-based intermediaries. Specifically, our contributions include:
  • Complex-valued holographic radiance fields. We present a new variant of 3DGS, utilizing complex-valued Gaussian primitives. Uniquely, our method is transformation-free and follows the geometry of 3D scenes in terms of the phase and amplitude, regardless of the viewpoints.
@@
  • Complex-valued differentiable wave-optics rasterizer. We enable efficient rendering and optimization of complex-valued holographic radiance fields by rasterizing optical waves emitted from a 3D scene.
We evaluate the image quality and speed of our method by comparing against existing Gaussian-based CGH approaches. Our codebase is publicly available at our project page (Zhan et al., 2025b).
@@ 2. Related Work
Our approach bridges recent advances in neural rendering with holography, enabling the efficient generation of holograms from novel views. We will first review novel-view synthesis methods and then examine existing wave optics rendering techniques that also attempt to simulate light phenomena beyond intensity. Finally, we discuss the traditional CGH methods in general, including point-based, polygon-based, and layer-based methods.
@@ 2.1. Novel-View Synthesis
Traditional novel-view synthesis methods primarily model intensity and appearance, limiting their utility for emerging holographic display technologies that require phase-aware representations. Early techniques such as image-based rendering (Chaurasia et al., 2013) and volumetric scene representations (Hedman et al., 2017; Riegler and Koltun, 2020) laid the groundwork for view-dependent appearance modeling. Building on these foundations, neural radiance fields (NeRF) (Mildenhall et al., 2021) advanced the field by learning continuous radiance fields that capture view-dependent color and opacity. This representation has since been extended to improve visual quality and computational efficiency (Barron et al., 2021, 2022, 2023; Chen et al., 2022; Rebain et al., 2021; Niemeyer et al., 2024; Duckworth et al., 2024).
Point-based representations have recently advanced with 3DGS (Kerbl et al., 2023), emerging as a promising alternative to NeRF. 3DGS offers a promising foundation for our work, as it encodes both geometry and appearance using oriented Gaussian primitives, but it is still fundamentally limited to intensity-based representations. Our work extends 3DGS to incorporate intrinsic amplitude and phase information, enabling it to model complex-valued holographic radiance fields that capture not only intensity but also interference and diffraction phenomena.
@@
The work by Kim  et al. (2023) extends NeRF beyond intensity-based representations to capture spectral and polarimetric properties of light, revealing richer material and structural characteristics of 3D scenes. Notably, spectral and polarization data can be supervised using ground-truth measurements obtained from specialized imaging devices. On the other hand, interference and diffraction phenomena arise from the superposition of phase and amplitude values. Interference and diffraction effects lack direct measurement data and must be inferred through computational means, as holographic cameras capable of capturing both amplitude and phase information are not yet commodity equipment (An et al., 2020). Moreover, supporting simulations of holography in rendering pipelines is a common theme in research. For instance, Steinberg  et al. (2022; 2024) introduced wave optics formulations for path tracing, enabling partial coherence and diffraction effects. Choi  et al. (2025) proposed Gaussian Wave Splatting (GWS) that transforms 2D Gaussian Splatting (Huang et al., 2024) scenes into holograms by deriving closed-form solutions for angular spectrum. Chen  et al. (2025) proposed a two-stage method that combines pretrained intensity-based 3DGS with U-Net (Ronneberger et al., 2015), enabling efficient hologram synthesis for novel views. In this paper, we propose to enable recalculation-free novel-view hologram synthesis, whereas Choi  et al. requires per-view wave re-computation and Chen  et al. requires per-view U-Net inference.
@@ 2.2. Eulerian-based Computer-Generated Holography
Holographic displays rely on CGH, which algorithmically synthesizes diffraction and interference patterns to reconstruct 3D scenes with near-accurate depth cues. Conceptually, traditional CGH methods follow a per-view Eulerian-based approach to wave propagation, where the coordinates on the hologram plane remain fixed, while the complex field varies when the relationship between the scene geometry and the viewpoint changes, requiring expensive recalculation. The computed holographic patterns are implemented using Spatial Light Modulator (SLM), which modulates the phase of incident coherent beams to generate designed image patterns. Through numerical wave propagation simulations (Kavaklı et al., 2022; Peng et al., 2020), CGH generates holograms leading to high spatio-angular resolution, positioning it as a promising emerging screen technology. The accuracy and computational complexity of CGH strongly depend on primitive representations, which are generally categorized into point-based, polygon-based, and layer-based methods.
Point-based CGH represents scenes using discrete points emitting spherical waves. The wave field at the hologram, $P$, is calculated by summing waves from N points,
@@
where $A_{n}$ denotes amplitude, $k$ is the wave number, and $r_{n}$ is the propagation distance from point $n$ to $P(x,y)$. Reflecting the Huygens-Fresnel principle, Eq. (1) is highly flexible and physically accurate. Maimone  et al. (2017) presented point-source-based CGH computation for holographic near-eye displays for virtual and augmented reality. Shi  et al. (2021) leveraged point clouds and triangle-ray intersection for occlusion handling in large-scale hologram datasets. However, the computational cost of point-based CGH increases significantly with point density, and occlusion handling introduces additional overhead.
@@
Polygon-based CGH balances efficiency and accuracy by representing surfaces using triangular or polygonal facets emitting coherent wave fields. This method naturally integrates shading models and intrinsic occlusion handling through visibility algorithms (e.g. back-face culling and affine transformations) (Zhang et al., 2018b; Matsushima and Nakahara, 2009). Mesh-based CGH includes topology information for precise occlusion handling (Ahrenberg et al., 2008; Yeom et al., 2022). While achieving realistic shading by combining wave optics with rendering pipelines, these methods suffer from increased algorithmic complexity and struggle with finely-detailed or non-polygonal structures.
@@
Layer-based CGH subdivides scenes into discrete depth planes, propagating complex fields from $P$ independently to each plane using operators like the Angular Spectrum Method (ASM) (Matsushima and Shimobaba, 2009; Zhang et al., 2020; Zheng et al., 2024). The $P(x,y)$ is optimized so the propagated results across all planes approximate the desired 3D scene. Layer-based CGH greatly accelerates computations through Fast Fourier Transform (FFT)-based propagation, making it suitable for faster rendering and is widely adopted in learned CGH methods (Shi et al., 2022; Choi et al., 2021; Peng et al., 2020; Zhan et al., 2024; Liu et al., 2023). Building on this acceleration, learned complex-valued CGH methods (Zhong et al., 2023; Zhang et al., 2025) generate phase-only holograms from complex-valued fields in a single forward pass. These approaches share the same application as our method for SLM-based holographic displays. In contrast, our method embeds complex-valued primitives directly into the 3D scene representation, eliminating the need for re-inference per viewpoint.
@@
While recent CGH advancements enable realistic defocus effects and continuous pupil accommodation (Kim et al., 2024; Kavaklı et al., 2023a), existing light field–based 3D and 4D CGH methods are typically optimized for limited viewing directions, with computed holograms losing the scene information under large viewpoint changes. We propose integrating strengths of both point-based and layer-based CGH methods, utilizing complex-valued 3D Gaussians as point-based primitives to encode scene geometry and complex-valued radiance. Unlike existing Eulerian-based CGH approaches, our method follows a Lagrangian-based formulation, treating amplitude and phase as intrinsic variables in scene representations. Compared with point clouds and polygon-meshes, Gaussians have learnable scale and rotation parameters, allowing them to capture anisotropic structures and fine-grained spatial variations effectively. This expressiveness enables both effective appearance modeling and geometrically faithful wave emitting distribution. Table 1 presents a comparison between our method and other Gaussian-based CGH approaches, demonstrating its ability to achieve fast inference, natural defocus blur, on-par image quality, and scene geometry-aware amplitude and phase representations across novel views.
Table 1. Comparison of Gaussian primitives-based hologram synthesis method. Inference refers to inference speed. Re refers to whether the method requires recalculation when geometric relationship between scene and observer changes. Type refers to Gaussian representation type. Quality refers to the image quality of reconstructed images from the holograms. Align refers to scene geometry-aware amplitude and phase representations across novel views. NDB refers to natural defocus blur.
@@
GWS (2025) Moderate Yes Intensity Good No No
@@
Figure 3. Without the need to query an intensity-based radiance field, our approach models a complex-valued holographic radiance field using Gaussian primitives with intrinsic amplitude and phase properties, enabling scene geometry-aware amplitude and phase modeling across viewpoints and efficient rendering through a differentiable multi-layer propagation pipeline.